# A Runtime-Adaptive Transformer Neural Network Accelerator on FPGAs

### v1

@@ 1. Introduction
Lu et al. (lu_hardware_2020, ) accelerated the attention mechanism and feedforward network separately, but did not implement the full transformer encoder. Ye et al. focused on accelerating only the attention mechanism using a reconfigurable systolic array for the transformer. Similarly, Zhang et al. (zhang_algorithm-hardware_2021, ) concentrated on accelerating the attention layer through hardware-software co-design. disabledisabletodo: disableIn this section you say ”some others papers” etc a lot and its hard to follow. Why not call a few papers out by name, like ”Ye et al. showed …” That would be easier to read. You also have a lot of ”transition words” like ”moreover, therefore, however”, to me, this means the section needs more structure and anyway to do that would be to call out individual papers form authors by name. (ye_accelerating_2023; zhang_algorithm-hardware_2021, ; lu_hardware_2020, ) neglecting feedforward networks and non-linear operations. Some other works accelerate the full transformer networks but their logic circuits go through the time-consuming synthesis steps for different models or they perform poorly on the same model with different configurations (FaFlA, ). They lack generality to support other variants.
@@ 2.1. Transformer Architecture
The output of MHA can be represented as Equation 1 & 2. The input sequence X is linearly mapped into $Q_{i},K_{i},V_{i}$ matrices using weights and biases. The parameter $d_{k}=d_{model}/h$ is the dimension of $Q_{i}$ and $K_{i}$. $d_{model}$ is a hyperparameter called embedding dimension, and $h$ is the number of heads. The mask operation filters out all values of illegal connections before the softmax. The parameter $d_{k}$ is set to 64 in both the transformer base model and the transformer large model.
(1) $\begin{split}Attention(Q_{i},K_{i},V_{i})=softmax\left(Mask\left(\frac{Q_{i}K^{T}_{i}}{\sqrt{d_{k}}}\right)\right)V_{i}\end{split}$
@@
(3) $\begin{split}{FFN}\_{ResBlock}(X)=&Layer\_Norm(X+\\ &ReLU(X\times W_{1}+b_{1})\times W_{2}+b_{2})\end{split}$
@@
The FFN ResBlock comprises a LN operation, residual addition, a ReLU activation, and two linear sublayers, as described in Equation 3, where W1, W2 are weights and b1, b2 are biases. The operations for layer normalization, softmax, GELU and RELU activation functions are described in equations 4, 5, 6, and 7 respectively, where x is the input vector (for a particular position in the sequence), $\mu$ is the mean of x, ${\sigma}^{2}$ is the variance of x, $\gamma$ and $\beta$ are learnable parameters, and $\epsilon$ is a small constant.
@@ 3. ADAPTOR’s Architecture
The core of the ADAPTOR is designed in C language on Vitis high-level synthesis (HLS) 2022.2.1 tool. C simulation confirms the algorithm’s correctness, while C/RTL co-simulation validates the functionality of the synthesized hardware. This section describes the HLS design technique that generates an optimized architecture utilizing most of the LUTs and DSPs in the processing modules, ensuring high parallelism of computation. The input vectors and weights are pre-stored in off-chip memory (DRAM/HBM) and transferred to the accelerator via AXI master interfaces. The accelerator receives control signals from the processor through an AXI-lite slave interface. There are loading units, computing modules and activation function units in the overall architecture, which are described below. Fig. 2 and 3 represent two main computing modules of ADAPTOR.
@@ 3.1. Load Inputs Unit
Inputs are loaded into an input BRAM from external memory. Input BRAM is a dual port BRAM implemented as a two-dimensional array of size ($SL\times d_{model}$) in HLS, where $SL$ is sequence__\__length. It is reused to store the outputs of each encoder/decoder layer so that it can send inputs to the subsequent layer. One Load_inputs unit loads data to intermediate input buffers of each attention head from external memory using Algorithm 1. These input buffers are declared as a two-dimensional array of size ($SL\times TS_{MHA}$). Therefore, tiling is applied along the column of the matrix, and they are replenished with new data ($\frac{d_{model}}{TS_{MHA}}$) times. Another Load_inputs unit loads data to the input buffers of the FFN1 module which are declared as a two-dimensional array of size ($SL\times TS_{FFN}$) using algorithm 3. The third Load_inputs unit loads data to the input buffers of the FFN2 module which are declared as a two-dimensional array of size ($SL\times(4\times TS_{FFN}$)) using algorithm 4. Different units are used because of different number of computations, tile size and array shapes.
3.2. Load Weights Unit
There are three Load_weight units. One unit loads weights to the weight buffers of each attention head according to demand from external memories. Another unit loads weights to the weight buffers of the feedforward network. The third one loads weights for the layer normalization module. Weight buffers of the attention module are declared as two-dimensional arrays of size ($\frac{d_{model}}{h}\times TS_{MHA}$) in HLS. $TS_{MHA}$ is tile_size for the attention module, which represents the dimension of the sub-matrices into which the larger weight matrices are divided. They are synthesized as dual-port BRAMs or LUTRAMs. Therefore, they are loaded with partial data from external memories for each tile at each iteration. Weight buffers of the feedforward network are declared as two-dimensional arrays of size ($TS_{FFN})\times(4\times TS_{FFN})$. TSFFN is tile_size for the FFN module which equals to $\frac{Embedding\_Dimension}{No.\_of\_Tiles\_FFN}$. (4 $\times$ TSFFN) equals to $\frac{Hidden\_Dimension}{No.\_of\_Tiles\_FFN}$. Thus, they are tiled along two dimensions (row and column of the matrix) and loaded iteratively for each tile in each dimension. They are also synthesized as dual-port BRAMs or LUTRAMs. The weight buffers for the layer normalization unit are declared as one-dimensional arrays of size $d_{model}$ in HLS. They are not tiled, and so, loaded at once, and they are converted into dual-port BRAMs or LUTRAMs after synthesis. Algorithm 2 describes the weight loading process.
@@ 3.3. Load Biases Unit
Algorithm 2 Load Weights for MHA
1:for $(i=1;i<=\frac{Embedding\ Dimension}{Number\ of\ Heads};i=i+1)$ do
2:     #pragma HLS pipeline off
3:     for $(j=1;j<=Tiles\_in\_MHA;j=j+1)$ do
4:         #pragma HLS pipeline II = 1
5:         $w_{q}[i][j]\leftarrow weights\_Q[index];$
6:         $w_{k}[i][j]\leftarrow weights\_K[index];$
7:         $w_{v}[i][j]\leftarrow weights\_V[index];$
8:         $index\leftarrow index+1;$
9:     end for
10:end for
@@
3.4. Activation Unit
3.5. Layer Normalization Unit
The mean and variance of the outputs from the attention layer and feedforward network layer are calculated in Layer Normalization (LN) unit according to Equation 4 which is implemented by algorithm 8 in HLS. The values are normalized using the mean and variance, and the results are multiplied by weights and then added with biases element-wise. The outputs from FFN1 and FFN3 go through the LN unit before further processing.
@@ 3.6. Attention Module
The architecture to accelerate the attention module is shown in Fig. 2. There are three main processing modules (PM) in it. They are denoted as ${QKV}_{PM}$, ${QK}_{PM}$ and ${SV}_{PM}$ according to the output they produce. The number of instances for these modules depends on the number of attention heads (h). Each module contains an array of processing elements (PE). A PE is comprised of a DSP48 performing multiplication and accumulation (MAC) operations. The number of PEs ($t$) depends on the unrolling factor of the inner loop and the initiation interval of the pipelined outer loop. The PE array’s data access pattern and computational requirements differ across modules. Therefore, they are defined separately as functions in HLS to produce distinct sets of PE arrays. This approach enables optimization of each RTL module separately. Input data and weights are stored in multiple BRAMs/LUTRAMs to enable parallel access.
In our architecture, each PE is independent, with its own local memory, control, and computing unit. The weights ($W_{q}$, $W_{k}$, $W_{v}$) for generating query (Q), key (K) and value (V) matrix are declared as separate two-dimensional arrays of size ($\frac{d_{model}}{h}\times TS_{MHA}$) in HLS. $TS_{MHA}$ is the tile size n the attention module. It represents the dimension of the sub-matrices into which the larger weight matrices are divided. The number of heads and tiles, and the array partition directive on HLS determine how the arrays will be partitioned to generate multiple two-port BRAMs. Due to the limited ports of BRAMs, array partitioning and data loading are efficiently managed to ensure that data required simultaneously by a DSP are stored in separate BRAMs. The $Q$, $K$, and $V$ matrices of size ($SL\times\frac{d_{model}}{h}$) are stored in intermediate buffers where $SL$ is the sequence length.
@@ 3.6.1. QKVPM module:
${QKV}_{PM}$ module generates the query, key, and value matrices. This module contains the $W_{q}$, $W_{k}$, $W_{v}$ buffers, and input ($X_{i}$) buffers from which data is accessed in parallel by parallel DSP units. The arrays used in this module are divided into subarrays using our tiling technique to fit into the BRAMs or LUTRAMs. The number of loop iterations of the ${QKV}_{PM}$ module depends on the tile size. There are a total of ($\frac{d_{model}}{TS_{MHA}}$) tiles or iterations. At each iteration, the $W_{q}$, $W_{k}$, $W_{v}$, and $X_{i}$ buffers are loaded with distinct data. Then the computations start in the PEs. Simultaneously, the biases for the $Q$, $K$, and $V$ matrices are loaded to registers from off-chip memory while the ${QKV}_{PM}$ module performs computations. Then they are added with $Q$, $K$, $V$ matrices. Algorithm 9 outlines the computations of this module where the 2nd loop (line 6) is pipelined causing the innermost loop (line 8) to be fully unrolled. This generates ($\frac{d_{model}}{TS_{MHA}}$) PEs.
@@ 3.7. Feedforward Network Module
There are three RTL modules for FFN to perform the operations of feedforward networks of different architectures. The definitions of the functions representing the modules have different dimensions of arrays for the inputs and outputs in HLS. These arrays are converted into BRAMs/LUTRAMs after synthesis. The number of computations inside each module is different, so each FFN has a separate function. They contain a different number of processing elements after synthesis because of different unrolling factors of the innermost loop. The weights are stored in a two-dimensional array ($W_{o}$) with dimensions ($\frac{d_{model}}{TS_{FFN}}\times\frac{4\times d_{model}}{TS_{FFN}}$) in HLS, where $TS_{FFN}$ represents the tile size in the FFN. Both $FFN1_{PM}$ and $FFN3_{PM}$ are followed by layer normalization (LN) modules.
Figure 2. Attention Module of ADAPTOR.
@@ 3.7.3. FFN3PM module:
3.8. Bias Add Unit
There are three Bias_add units for adding biases to the query ($Q$), key ($K$), and value ($V$) matrices and the outputs of the three feedforward networks. Three units are used because the HLS function definition corresponding to a unit takes arrays of different dimensions as inputs and generates arrays of different dimensions as outputs. One of the units used after feedforward networks contains the ReLU activation function. Algorithm 15, 16 & 17 are the operations of the three units.
@@ 3.8. Bias Add Unit
10:              $S_{q}\leftarrow S_{q}+x[i][j]\times w_{k}[k][j];$
11:              $S_{q}\leftarrow S_{q}+x[i][j]\times w_{v}[k][j];$
@@
1:for $(i=1;i<=Sequence\ Length;i++)$ do
2:     #pragma HLS pipeline off
3:     for $(j=1;j<=Hidden\ Dimension;j++)$ do
4:         #pragma HLS pipeline II = 1
5:         $FF_{out}[i][j]\leftarrow FF_{out}[i][j]+bias_{FFN}[j];$
6:         $FF_{out}[i][j]\leftarrow relu(FF_{out}[i][j]);$
7:     end for
8:end for
Algorithm 17 Bias add unit 3
3.9. Tiling Technique
Figure 3. Feedforward Network Module of ADAPTOR.
@@ 3.9. Tiling Technique
3.10. Tile Size Determination
In ADAPTOR, the programmable parameters can be adjusted at runtime, whereas the tile size must be set before synthesis, as it cannot be modified without re-synthesizing the entire hardware. The graph in Fig. 5(a) & (b) illustrates how variations in $TS_{MHA}$ and $TS_{FFN}$ impact system frequency (MHz) and latency (normalized to the minimum value), respectively. The number of tiles in MHA ($\frac{d_{model}}{TS_{MHA}}$) was varied from 6 to 48, for each FFN tile count ($\frac{d_{model}}{TS_{FFN}}$) which ranged from 2 to 6. The plots indicate that the optimal configuration for achieving the highest frequency and the lowest latency was 12 tiles in MHA and 6 tiles in FFN. This combination would achieve a maximum frequency of 200 MHz. Moreover, experiments showed that $TS_{MHA}$ of 64 and $TS_{FFN}$ of 128 are optimal for HLS, allowing for efficient array partitioning within a reasonable compilation time (approximately 36 hours) for a state-of-the-art (SOTA) transformer.

### v2

@@ 1. Introduction
Lu et al. (lu_hardware_2020, ) accelerated the attention mechanism and feedforward network separately, but did not implement the full transformer encoder. Ye et al. (ye_accelerating_2023, ) focused on accelerating only the attention mechanism using a reconfigurable systolic array for the transformer. Similarly, Zhang et al. (zhang_algorithm-hardware_2021, ) concentrated on accelerating the attention layer through hardware-software co-design. disabledisabletodo: disableIn this section you say ”some others papers” etc a lot and its hard to follow. Why not call a few papers out by name, like ”Ye et al. showed …” That would be easier to read. You also have a lot of ”transition words” like ”moreover, therefore, however”, to me, this means the section needs more structure and anyway to do that would be to call out individual papers form authors by name. (ye_accelerating_2023, ; zhang_algorithm-hardware_2021, ; lu_hardware_2020, ) neglecting feedforward networks and non-linear operations. Some other works accelerate the full transformer networks but their logic circuits go through the time-consuming synthesis steps for different models or they perform poorly on the same model with different configurations (FaFlA, ). They lack generality to support other variants.
@@ 2.1. Transformer Architecture
The output of MHA can be represented as Equation 1 & 2. The input sequence X is linearly mapped into $Q_{i},K_{i},V_{i}$ matrices using weights and biases. The parameter $d_{k}=d_{model}/h$ is the dimension of $Q_{i}$ and $K_{i}$. $d_{model}$ is a hyperparameter called embedding dimension, and $h$ is the number of heads.
(1) $\begin{split}Attention(Q_{i},K_{i},V_{i})={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}}softmax\left(\frac{Q_{i}K^{T}_{i}}{\sqrt{d_{k}}}\right)V_{i}\end{split}$
@@
(3) $\begin{split}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}}FFN(X)=&Layer\_Norm(X+\\ &ReLU(X\times W_{1}+b_{1})\times W_{2}+b_{2})\end{split}$
@@
The FFN comprises a LN operation, residual addition, a ReLU activation, and two linear sublayers, as described in Equation 3, where W1, W2 are weights and b1, b2 are biases. The operations for layer normalization, softmax, GELU and RELU activation functions are described in equations 4, 5, 6, and 7 respectively, where X is the input vector (for a particular position in the sequence), $\mu$ is the mean of X, ${\sigma}^{2}$ is the variance of X, $\gamma$ and $\beta$ are learnable parameters, and $\epsilon$ is a small constant.
@@ 3. ADAPTOR’s Architecture
The core of the ADAPTOR is designed in C language on Vitis high-level synthesis (HLS) 2022.2.1 tool. C simulation confirms the algorithm’s correctness, while C/RTL co-simulation validates the functionality of the synthesized hardware. This section describes the HLS design technique that generates an optimized architecture utilizing most of the LUTs and DSPs in the processing modules, ensuring high parallelism of computation. There are loading units, computing modules, and activation function units in the overall architecture, which are described below. Fig. 2 and 3 represent two main computing modules of ADAPTOR.
3.1. Load Weights Unit
There are three Load_weight units. One unit loads weights to the weight BRAMs (Fig. 2) of each attention head according to demand from external memories. Another unit loads weights to the weight BRAMs (Fig. 3) of the feedforward network. The third one loads weights for the layer normalization module. Weight BRAMs of the attention module are declared as two-dimensional arrays of size ($\frac{d_{model}}{h}\times TS_{MHA}$) in HLS. $TS_{MHA}$ is tile_size for the attention module, which represents the dimension of the sub-matrices into which the larger weight matrices are divided. They are synthesized as dual-port BRAMs or LUTRAMs. Therefore, they are loaded with partial data from external memories for each tile at each iteration. Weight BRAMs of the feedforward network are declared as two-dimensional arrays of size ($TS_{FFN})\times(4\times TS_{FFN})$. TSFFN is tile_size for the FFN module which equals to $\frac{Embedding\_Dimension}{No.\_of\_Tiles\_FFN}$. (4 $\times$ TSFFN) equals to $\frac{Hidden\_Dimension}{No.\_of\_Tiles\_FFN}$. Thus, they are tiled along two dimensions (row and column of the matrix) and loaded iteratively for each tile in each dimension. They are also synthesized as dual-port BRAMs or LUTRAMs. The weight BRAMs for the layer normalization unit are declared as one-dimensional arrays of size $d_{model}$ in HLS. They are not tiled, and so, loaded at once, and they are converted into dual-port BRAMs or LUTRAMs after synthesis. Algorithm 1 describes the weight loading process.
@@ 3.2. Load Inputs Unit
Inputs are loaded into an input BRAM from external memory. It is a dual port BRAM implemented as a two-dimensional array of size ($SL\times d_{model}$) in HLS, where $SL$ is sequence__\__length. It is reused to store the outputs of each encoder/decoder layer so that it can send inputs to the subsequent layer. One Load_inputs unit loads data to the intermediate Input BRAMs (Fig. 2) of each attention head from external memory using algorithm 2. These BRAMs are declared as a two-dimensional array of size ($SL\times TS_{MHA}$). Therefore, tiling is applied along the column of the matrix, and they are replenished with new data ($\frac{d_{model}}{TS_{MHA}}$) times. Another Load_inputs unit loads data to the Score BRAMs (Fig. 3) of the FFN1 module which are declared as a two-dimensional array of size ($SL\times TS_{FFN}$) using algorithm 3. The third Load_inputs unit loads data to the LN1 Buffers (Fig. 3) of the FFN2 module which are declared as a two-dimensional array of size ($SL\times(4\times TS_{FFN}$)) using algorithm 4. Different units are used because of different numbers of computations, tile sizes, and array shapes.
@@ 3.3. Load Biases Unit
3.4. Activation Unit
3.5. Layer Normalization Unit
The mean and variance of the outputs from the attention layer and feedforward network layer are calculated in Layer Normalization (LN) unit according to Equation 4 which is implemented by algorithm 8 in HLS. The values are normalized using the mean and variance, and the results are multiplied by weights and then added with biases element-wise. The outputs from FFN1 and FFN3 go through the LN unit before further processing (Fig. 3).
Algorithm 1 Load Weights for MHA
1:for $(i=1;i<=\frac{Embedding\ Dimension}{Number\ of\ Heads};i=i+1)$ do
2:     #pragma HLS pipeline off
3:     for $(j=1;j<=Tiles\_in\_MHA;j=j+1)$ do
4:         #pragma HLS pipeline II = 1
5:         $W_{Q}[i][j]\leftarrow weights\_Q[index];$
6:         $W_{K}[i][j]\leftarrow weights\_K[index];$
7:         $W_{V}[i][j]\leftarrow weights\_V[index];$
8:         $index\leftarrow index+1;$
9:     end for
10:end for
@@ 3.5. Layer Normalization Unit
Figure 2. Attention Module of ADAPTOR.
@@ 3.6. Attention Module
The architecture to accelerate the attention module is shown in Fig. 2. There are three main processing modules (PM) in it. They are denoted as ${QKV}_{PM}$, ${QK}_{PM}$ and ${SV}_{PM}$ according to the output they produce. The number of instances for these modules is equal to the number of attention heads (h). Each module contains an array of processing elements (PE). A PE is comprised of a DSP48 that performs multiplication and accumulation (MAC) operations. The PE array’s data access pattern and computational requirements differ across modules. Therefore, they are defined separately as functions in HLS to produce distinct sets of PE arrays. This approach enables optimization of each RTL module separately. Input data and weights are stored in multiple BRAMs/LUTRAMs to enable parallel access.
In our architecture, each PE is independent, with its own local memory, control, and computing unit. The weights ($W_{q}$, $W_{k}$, $W_{v}$) for generating query (Q), key (K) and value (V) matrix are declared as separate two-dimensional arrays of size ($\frac{d_{model}}{h}\times TS_{MHA}$) in HLS. The number of heads and tiles, and the array partition directive on HLS determine how the arrays will be partitioned to generate multiple two-port BRAMs. Due to the limited ports of BRAMs, array partitioning and data loading are efficiently managed to ensure that data required simultaneously by a DSP are stored in separate BRAMs. The $Q$, $K$, and $V$ matrices of size ($SL\times\frac{d_{model}}{h}$) are stored in intermediate buffers.
@@ 3.6.1. QKVPM module:
${QKV}_{PM}$ module generates the query, key, and value matrices. This module contains the $W_{Q}$, $W_{K}$, $W_{V}$ BRAMs, and input ($X_{i}$) BRAMs from which data is accessed in parallel by parallel DSP units. The arrays used in this module are divided into subarrays using our tiling technique to fit into the BRAMs or LUTRAMs. The number of times the ${QKV}_{PM}$ module is accessed depends on the tile size. There are a total of ($\frac{d_{model}}{TS_{MHA}}$) tiles or iterations. At each iteration, the $W_{Q}$, $W_{K}$, $W_{V}$, and $X_{i}$ buffers are loaded with distinct data. Then the computations start in the PEs. Simultaneously, the biases for the $Q$, $K$, and $V$ matrices are loaded to registers from off-chip memory while the ${QKV}_{PM}$ module performs computations. Then they are added with $Q$, $K$, $V$ matrices. Algorithm 9 outlines the computations of this module where the 2nd loop (line 6) is pipelined causing the innermost loop (line 8) to be fully unrolled. This generates ($\frac{d_{model}}{TS_{MHA}}$) PEs.
@@ 3.7. Feedforward Network Module
The architecture to accelerate the feedforward network module is shown in Fig. 3. There are three RTL modules for FFN to perform the operations of feedforward networks of different architectures. The definitions of the functions representing the modules have different dimensions of arrays for the inputs and outputs in HLS. These arrays are converted into BRAMs/LUTRAMs after synthesis. The number of computations inside each module is different, so each FFN has a separate function. They contain a different number of processing elements after synthesis because of different unrolling factors of the innermost loop. The weights are stored in a two-dimensional array ($W_{o}$) with dimensions ($\frac{d_{model}}{TS_{FFN}}\times\frac{4\times d_{model}}{TS_{FFN}}$) in HLS, where $TS_{FFN}$ represents the tile size in the FFN. Both $FFN1_{PM}$ and $FFN3_{PM}$ are followed by layer normalization (LN) modules.
@@ 3.7.3. FFN3PM module:
Figure 3. Feedforward Network Module of ADAPTOR.
3.8. Bias Add Unit
There are three Bias_add units for adding biases to the query ($Q$), key ($K$), and value ($V$) matrices and the outputs of the three feedforward networks. Three units are used because the HLS function definition corresponding to a unit takes arrays of different dimensions as inputs and generates arrays of different dimensions as outputs. One of the units used after feedforward networks contains the ReLU activation function. Algorithm 16, 17 & 13 are the operations of the three units.
3.9. Tiling Technique
@@
3.10. Tile Size Determination
In ADAPTOR, the programmable parameters can be adjusted at runtime, whereas the tile size must be set before synthesis, as it cannot be modified without re-synthesizing the entire hardware. The graph in Fig. 5(a) & (b) illustrates how variations in $TS_{MHA}$ and $TS_{FFN}$ impact system frequency (MHz) and latency (normalized to the minimum value), respectively. The number of tiles in MHA ($\frac{d_{model}}{TS_{MHA}}$) was varied from 6 to 48, for each FFN tile count ($\frac{d_{model}}{TS_{FFN}}$) which ranged from 2 to 6. The plots indicate that the optimal configuration for achieving the highest frequency and the lowest latency was 24 tiles in MHA and 6 tiles in FFN. This combination would achieve a maximum frequency of 200 MHz.
@@ 3.10. Tile Size Determination
10:              $S_{k}\leftarrow S_{k}+X[i][j]\times W_{K}[k][j];$
11:              $S_{v}\leftarrow S_{v}+X[i][j]\times W_{V}[k][j];$
@@
1:for $(i=1;i<=Sequence\ Length;i++)$ do
2:     #pragma HLS pipeline off
3:     for $(j=1;j<=Hidden\ Dimension;j++)$ do
4:         #pragma HLS pipeline II = 1
5:         $FF_{out}[i][j]\leftarrow FF_{out}[i][j]+bias_{FFN}[j];$
6:         $FF_{out}[i][j]\leftarrow relu(FF_{out}[i][j]);$
7:     end for
8:end for
Algorithm 13 Bias add unit 3
@@
3.11. Software Control